OpenStudy (anonymous):

The sum of 6 times Petra's age and 8 times Kathy's age is 162. Kathy is 1 year more than twice as old as Petra is. Find each of their ages

OpenStudy (anonymous):

Let Petra's age be x Let Kathy's age be y 6x+8y=162 <---------1 equation y=2x+1 6x + 8(2x+1)=162 6x+16x+8=162 22x=154 x=7 Substituting x in y=2x+1 y=2(7)+1 y=14+1 y=15